# Attention vs Choice in Welfare Take-up: What Works for WIC?

### v3

@@ Abstract
We decompose the causes of incomplete take-up of welfare program benefits into two latent mechanisms: inattention, where households do not consider program participation, and active choice, where households consider participation but find it not worthwhile. To capture these two mechanisms, we model households’ take-up decision as a two-stage process: attention followed by choice. Our model allows covariates to influence both stages simultaneously, creating an identification challenge in quantifying covariates’ impact on either stage separately. Exploiting two institutional structures of welfare programs: full attention among current participants and the recertification process, we prove that attention and choice parameters are constructively identifiable. Applied to National Longitudinal Survey of Youth 1997 (NLSY97) data on the Special Supplemental Nutrition Program for Women, Infants, and Children (WIC), our model reveals substantial household-level heterogeneity in both attention and choice probabilities. Furthermore, counterfactual simulations predict that choice-nudging policies outperform attention-boosting policies. We test this prediction using data from a WIC pilot program in Vermont that sent choice-nudging and attention-boosting text messages to different households. Consistent with the counterfactual prediction, choice-nudging messages increased retention by 8–10%, while attention-boosting messages had statistically insignificant effects. Our results highlight the critical role of choice nudges in improving program take-up. (JEL: I38, H75, D12, D91, C33, C54)
@@ 1 Introduction
Despite extensive evidence of welfare programs’ positive impacts on participants’ well-being (Banerjee et al., 2024), participation remains stubbornly low (Ko and Moffitt, 2024). To address this gap between potential and realized impact, policymakers have long pursued two broad strategies: attention-raising policies, which increase the probability that eligible households consider participation, and choice-inducing policies, which increase the probability that households choose participation conditional on consideration, see examples in Bhargava and Manoli (2015); Ganong and Liebman (2018). [^1] Bhargava and Manoli (2015) collaborates with IRS to send out attention-raising mails that inform recipients of the eligibility rules of Earned Income Tax Credit; Ganong and Liebman (2018) documents six state-level choice-inducing policies, which reduce the hassle cost of sign-up and recertification. This policy landscape motivates two fundamental questions: What drives down the welfare program participation rate among eligible households, inattention or a low choice tendency? What kind of policies, attention-raising or choice-inducing, promote take-up more effectively?
While extensive research has documented incomplete welfare take-up and evaluated the effectiveness of various interventions (e.g., Currie (2004)), the central question of whether low participation stems primarily from inattention or from active choice remains unresolved. The closest attempts come from innovative experimental studies that compare treatments designed for either raising attention or inducing choice (Bhargava and Manoli, 2015; Finkelstein and Notowidigdo, 2019). [^2] Finkelstein and Notowidigdo (2019) has one treatment that targets these two mechanisms simultaneously. These studies provide suggestive evidence on the relative importance of attention and choice, but their main focus is on estimating policy treatment effects rather than disentangling the contribution of each mechanism in the take-up decision. Our paper addresses this gap by developing a two-stage framework that explicitly quantifies the contributions of inattention and low choice tendency to welfare program nonparticipation.
@@
We model households’ welfare take-up decision as a two-stage process where households must first pay attention before making a choice about program participation. At the attention stage, households that are already enrolled in the last period are fully attentive, whereas those that are not enrolled last period are stochastically attentive. The probability of paying attention for stochastically attentive households is a function of observed household characteristics. At the choice stage, households weigh the benefit of participation against its costs. Costs include both the hassle cost for sign-up or recertification and the usage cost associated with using the program benefit, e.g., stigma. We allow all covariates except for the recertification status to enter simultaneously into the attention and choice stages. Recertification status enters into the choice stage exclusively. This setup creates an identification challenge: when a covariate correlates with take-up, it is not immediately clear whether it operates through attention, choice, or both.
@@
We exploit two institutional structures of welfare programs to devise a novel strategy that separately identifies attention and choice parameters using standard panel data.
@@
  • Full attention among participants: Many welfare programs require ongoing engagement with the program, e.g., using the food voucher or the benefit card provided by the program when buying grocery. Therefore, current program participants are fully attentive in the next period due to their ongoing engagement, while non-participants are only stochastically attentive. This institutional structure allows us to decompose the information in welfare participation panel data along two distinct dimensions. The first dimension is the probability of staying in the program. This probability is a function of choice parameters only, since current participants are fully attentive and proceed directly to the choice stage. We show that this first dimension identifies all choice parameters except for the hassle parameters. [^3] The choice stage design partitions the choice parameters into three types: benefit parameters, hassle parameters, and usage cost parameters. The second dimension is the probability of starting participation, which depends on both attention and choice. By comparing these two dimensions, we identify the attention parameters.
@@
  • Recertification: Welfare programs are means-tested, involving a periodic eligibility recertification process. Households required to recertify their eligibility need to pay a hassle cost to continue participation. All else equal, comparing the continuation behavior of households in recertification periods to those not in recertification periods provides clean variation that identifies the hassle parameters. For our empirical application to WIC, the recertification process is administratively identical to the initial sign-up process, so our model assumes identical hassle for these two processes. In a later extension, we show that when recertification hassle and sign-up hassle differ, a program complexity measure proposed by Kleven and Kopczuk (2011) is sufficient to identify hassle costs.
We apply our model to study participation in the Special Supplemental Nutrition Program for Women, Infants, and Children (WIC) using data from the National Longitudinal Survey of Youth 1997 (NLSY97). In our sample, each participant receives, on average, more than \$100 in monthly benefits. Despite the sizable benefit, the take-up rate among eligible households hovers around 50% over our sample period from January 2000 to December 2009, a figure that aligns with the national average from the US Department of Agriculture. This persistent gap, along with the decade of policy efforts aimed at closing it (Davis, Leavitt and Chau, 2022), motivates our empirical application.
Our model finds substantial household-level heterogeneity in both attention and choice behaviors. We preview some of the estimation results of our most preferred specification. These results are numerically similar across all specifications in the paper and align with many existing findings.
  1. Infant: Having an infant is a strong attention trigger that prompts households to consider WIC participation. This finding is consistent with the existing long-standing literature on WIC’s benefit structure and take-up pattern, see Bitler, Currie and Scholz (2003); Tiehen (2014); Carlson and Neuberger (2021). Our model quantifies that a baseline household has an attention probability of 20.0%. [^4] A baseline household is defined: one infant, no preschooler, in a county where the local accessibility (defined in Section 4.6 is 15, high school education, non-Black, non-Hispanic, does not participate in SNAP, with an annual income of \$6600 per person in the year 2009. Our model involves a random effect. We set the random effect to 0 for the baseline household. All feature values are set either as the median or the mode. All else equal, without an infant, the baseline household’s attention probability sharply drops to 7.4%.
  2. Hassle cost: Our model estimates a substantial sign-up and recertification hassle cost, a common phenomenon in social assistance programs (Kleven and Kopczuk, 2011; Shepard and Wagner, 2025). A baseline household perceives the hassle of sign-up/recertification as large as \$97, close to two-thirds of one month’s worth of benefit for such a household (\$150). Similar to Rossin-Slater (2013), our model finds that local accessibility significantly reduces hassle cost. One unit increase in local accessibility, a measure defined as the log product of different types of facilities available in the residence county and mostly falls between the range of 5 to 20, lowers hassle cost by about \$8.
  3. Racial disparity: Our model estimates substantial racial disparities in both attention and choice for WIC. Black households tend to participate less than baseline households, whereas Hispanic households tend to participate more. [^5] Race of the household is defined as the race of the NLSY respondents. This finding aligns with our raw data and Liu and Liu (2016). Our model allows us to further attribute this racial disparity to the latent attention mechanism. According to our estimates, all else equal, a black household is 5.1 percentage points (pp) less likely to be attentive than a baseline household, whereas a Hispanic household is 4.3pp more likely to be attentive than a baseline household.
  4. SNAP: Consistent with Bitler, Currie and Scholz (2003), we find that WIC participation is strongly associated with SNAP participation. All else equal, our estimates that SNAP participation is associated with an 8.2pp increase in attention probability for a baseline household.
@@
  5. Unobserved heterogeneity: We find substantial household-level unobserved heterogeneity. This finding suggests that the incorporation of unobserved heterogeneity into both stages of the inattention model, as practiced by Heiss et al. (2021); Agarwal and Somaini (2025), can be crucial for counterfactual analysis based on such models.
@@
To evaluate potential policies, we use our preferred specification to simulate two types of counterfactuals: a population-wide perfect intervention and a realistic intervention that reflects feasible policy options.
@@
Under the population-wide perfect intervention counterfactual, we simulate two extreme policies: one raises the attention probability for all households to 100%, and the other induces a 100% choice probability for all households. Forcing attention increases the take-up rate by 41pp, whereas forcing choice increases it by 8pp. This result indicates that at the population level, inattention is the largest barrier to WIC participation.
@@
Given the infeasibility of population-wide perfect interventions, we simulate more realistic policies that reflect real-world constraints. WIC policymakers often focus retention efforts on currently-enrolled households due to budget limitations. We therefore design two types of more realistic interventions characterized as follows. One, an attention boosting intervention increases the attention probability of the household that recently exited the program to 100%, giving the household one more chance to choose participation. Two, a choice-nudging intervention increases the current participants’ next period choice probability by a small amount. Strikingly, we find that a choice-nudging intervention that increases participating households’ choice probability every month by a mere 0.45pp is as effective at raising overall take-up as an attention-boosting intervention that ensures full attention—but only for households in the month after they exit the program. The comparison suggests that under the constraints that policymakers can only intervene current participants, choice nudge is likely a more effective policy.
@@
Based on our realistic intervention counterfactual results, we hypothesize that a choice-nudging policy is likely to be more effective than an attention-boosting policy. We test this hypothesis using the WIC2Five experimental data. In 2016 and 2017, the Vermont WIC policymakers experimented with two types of text messages in different counties, both of which aimed at increasing the WIC retention rate. The first type of message was attention-boosting. Such messages reminded households that they missed an appointment, essentially asking the households to reconsider their choice. The second type of message was choice-nudging. Each month, the WIC policymakers sent out educational messages to inform households of the benefits of continuing WIC participation. The WIC2Five report concludes that the choice-nudging intervention “did not meet our goal of increasing child retention” by comparing the levels from July 2015 to July 2017. We reanalyze the retention rates data provided by the WIC2Five report by comparing the differences between treated and control sites. Using difference-in-differences (DiD), event study, and permutation test designs, our analysis yields opposing results from the WIC2Five report. Sites that received attention-boosting text messages saw virtually no change in their retention rate; sites that received choice-nudging messages witnessed a significant 8-10pp increase in retention. The results corroborate our counterfactual analysis: a choice-nudging policy is likely to be more effective than an attention-boosting policy when it comes to promoting WIC take-up.
@@
We contribute to three sets of literature. First, the existing works in the welfare take-up literature focus on estimating the causal relationship between observed covariates and the take-up rate (Rossin-Slater, 2013; Chareyron, Domingues and Lieno-Gaillardon, 2021) or the treatment effect of various policy interventions (Friedrichsen, König and Schmacker, 2018; Shepard and Wagner, 2025). Our work complements this literature by providing a structural framework that disentangles the attention and choice mechanisms behind these causal findings. Second, we propose new identifying conditions that have not been considered by existing inattention model literature (Hortaçsu, Madanizadeh and Puller, 2017; Heiss et al., 2021; Barseghyan, Molinari and Thirkettle, 2021; Abaluck and Adams-Prassl, 2021; Einav, Klopack and Mahoney, 2025; Agarwal and Somaini, 2025). Our identifying conditions are tied to the two institutional structures of welfare programs, making our model applicable to many other welfare take-up contexts. We further compare our methodological contribution to the existing inattention model literature in Appendix˜A. Third, we contribute to the long-standing empirical research on WIC (Bitler, Currie and Scholz, 2003; Rossin-Slater, 2013; Tiehen, 2014; Carlson and Neuberger, 2021; Davis, Leavitt and Chau, 2022). Using different methodologies and datasets, we present two pieces of new evidence that a choice-nudging intervention is likely to be more effective than an attention-boosting intervention in terms of promoting WIC take-up. Further experimental studies to test this hypothesis can be an interesting future research direction.
@@ 2.2 Identification
Following the consideration set literature (Abaluck and Adams-Prassl, 2021; Barseghyan, Molinari and Thirkettle, 2021; Agarwal and Somaini, 2025), we assume that we observe the take-up probability, $P(D_{it}=1\mid D_{it-1},X_{it},Z_{it})$, of a household for any given $\{D_{it-1},X_{it},Z_{it}\}\in\{0,1\}\times\mathcal{X}\times\{0,1\}$. Given the sequential nature of attention and choice, we decompose the observed single-period take-up probability into attention and choice probabilities under Assumption˜2.1 imposed on $\{\mathcal{E}_{it},\Xi_{it}\}$. We will explicitly relax this assumption in the full model.
@@ 3.1 Model setup
Assumption˜2.1 required that the attention utility shock $\mathcal{E}_{it}$ and the choice utility shock $\Xi_{it}$ are i.i.d. over time and mutually independent. The full model relaxes each of these restrictions by allowing for persistent, unobserved household heterogeneity captured by a random effect $Q_{i}$. Intuitively, $Q_{i}$ represents latent time-invariant household traits such as thriftiness (or some part of thriftiness). One component of thriftiness may be explained by observed covariates entering the utility functions, while the orthogonal residual component is unobserved and summarized by $Q_{i}$. In one of the model extensions, we discuss how to relax the orthogonality assumption.
@@ 3.3 Identification
Under Assumption˜3.1, any $\beta^{\omega}\in\beta^{\upsilon}\cup\beta^{\chi}$ are constructively identified as
@@
Under Assumption˜3.1, any $\beta^{\omega}\in\beta^{\kappa}$ are constructively identified as

### v5

@@ Abstract
Incomplete take-up of welfare benefits remains a major policy puzzle. This paper decomposes the causes of incomplete welfare take-up into two mechanisms: inattention, where households do not consider program participation, and active choice, where households consider participation but find it not worthwhile. To capture these two mechanisms, we model households’ take-up decision as a two-stage process: attention followed by choice. Applied to NLSY97 data on the Special Supplemental Nutrition Program for Women, Infants, and Children (WIC), our model reveals substantial household-level heterogeneity in both attention and choice probabilities. Furthermore, counterfactual simulations predict that choice-nudging policies outperform attention-boosting policies. We test this prediction using data from the WIC2Five pilot program that sent choice-nudging and attention-boosting text messages to different households. Consistent with the counterfactual prediction, choice-nudging messages increased retention much more effectively than attention-boosting messages. (JEL: I38, H75, D12, D91, C33, C54)
@@ 1 Introduction
Welfare programs significantly improve participants’ well-being (banerjee2024social). Yet it remains a policy puzzle why participation rates of many programs remain stubbornly low (ko2024take). One prominent example is the Special Supplemental Nutrition Program for Women, Infants, and Children (WIC). kreider2016identifying; sonchak2016impact show that WIC successfully reduces childhood food insecurity and increases important health measures such as birth weight. Despite these positive impacts, the take-up rate of WIC has hovered around 50% over the past two decades, leaving millions of eligible households out of the program and more than a billion dollars worth of benefits unclaimed per year.
To address the gap between high program effectiveness and low take-up rates, policymakers have long pursued two broad strategies: attention-raising policies and choice-inducing policies. Attention-raising policies increase the probability that eligible households consider participation. They typically involve running informational campaigns that promote program awareness or informing selected households about likely eligibility. In contrast, choice-inducing policies increase the probability that households choose participation conditional on consideration. They typically involve improving the perceived value of the program benefits or reducing the hassle involved in the sign-up and eligibility recertification processes. For example, in 2016, the Vermont WIC office sent out attention-boosting and choice-nudging text messages to WIC participants at different sites. Attention-boosting messages remind recently dropped-out households that they might still be eligible for WIC. Choice-nudging messages inform participating households that continued enrollment in WIC is beneficial for the future of the children, improving participants’ perceived value of the program benefits. [^1] It will become clear why these two interventions are attention-boosting and choice-nudging, respectively, once the two-stage inattention model is introduced. Both interventions aim at increasing participation by improving program retention. i.e., the intensive margin of program participation. [^2] See examples of other attention-raising and choice-inducing policy interventions for various welfare programs in bhargava2015psychological; ganong2018decline; finkelstein2019take. These interventions increase participation by increasing new program recruitment. i.e., the extensive margin of program participation. Our two-stage inattention model can accommodate policy counterfactuals for both policies that increase program retention and policies that increase new recruitment.
The dichotomy of attention-raising versus choice-inducing policies motivates two empirical questions: What is the primary driver for incomplete take-up, inattention or a low choice tendency? What kind of policies, attention-raising or choice-inducing, promote take-up more effectively? We study these two questions in the context of WIC.
@@
To answer the first question, we develop a two-stage inattention model that explicitly quantifies the contributions of inattention and low choice tendency to WIC nonparticipation. In our two-stage inattention model, households must first pay attention before making a choice about program participation. For a household to participate in WIC, it must pay attention at the attention stage and think that it is worthwhile to participate in WIC at the choice stage. We allow almost all covariates, including the lagged WIC take-up decision and household-level unobserved heterogeneity, to enter simultaneously into the attention and choice stages. This setup creates an identification challenge: when a covariate correlates with take-up, it is not immediately clear whether it operates through attention, choice, or both.
@@
We exploit two institutional structures of welfare programs to devise a novel strategy that separately identifies attention and choice parameters using standard panel data: full attention among participants and the presence of a recertification process. These two institutional structures are common in many welfare programs, making our model generalizable.
@@
Full attention among participants: Many welfare programs require ongoing engagement with the program. To use the WIC benefit package, enrollees actively redeem food vouchers or use the benefit card provided by the program when buying groceries. Therefore, current program participants are fully attentive in the next period due to their ongoing engagement, while non-participants are only stochastically attentive. This institutional structure allows us to decompose the information in welfare participation panel data along two distinct dimensions. The first dimension is the probability of staying in the program. This probability is a function of choice parameters only, since current participants are fully attentive and proceed directly to the choice stage. We show that this first dimension identifies all choice parameters except for the hassle parameters. [^3] The choice stage design partitions the choice parameters into three types: benefit parameters, hassle parameters, and usage cost parameters. The second dimension is the probability of starting participation, which depends on both attention and choice. By comparing these two dimensions, we identify the attention parameters.
@@
Recertification: Welfare programs are means-tested, involving a periodic eligibility recertification process. Households required to recertify their eligibility need to pay a hassle cost to continue participation. All else equal, comparing the continuation behavior of households in recertification periods to those not in recertification periods provides clean variation that identifies the hassle parameters.
@@
We apply our model to study WIC participation using data from the National Longitudinal Survey of Youth 1997 (NLSY97). Our model finds substantial household-level heterogeneity in both attention and choice behaviors. Notably, we find substantial household-level unobserved heterogeneity. This finding suggests that the incorporation of unobserved heterogeneity into both stages of the inattention model, as practiced by heiss2021inattention; agarwal2025demand, can be crucial for counterfactual analysis based on such models.
@@
To answer the second question, we simulate various attention-raising and choice-inducing policies using our model estimates. We use our preferred specification to simulate two types of counterfactuals: a population-wide perfect intervention and a realistic intervention that reflects feasible policy options.
@@
Population-wide perfect intervention result indicates that at the population level, inattention is the largest barrier to WIC participation. Under the population-wide perfect intervention counterfactual, we simulate two extreme policies: one raises the attention probability for all households to 100% during all periods, and the other induces a 100% choice probability for all households during all periods. Forcing attention increases the take-up rate by 42pp, whereas forcing choice increases it by 8pp.
@@
Given the infeasibility of population-wide perfect interventions, we simulate more realistic policies that reflect real-world constraints. WIC policymakers often focus retention efforts on currently-enrolled households due to budget limitations. We therefore design two policy counterfactuals that resemble the likely impact of the attention-boosting and choice-nudging text message interventions in the 2016 Vermont pilot program. One, an attention boosting intervention increases the attention probability of the household that recently exited the program to 100%, giving the household one more chance to choose participation. Two, a choice-nudging intervention increases the current participants’ next period choice probability by a small amount.
Strikingly, we find that a choice-nudging intervention that increases participating households’ choice probability every month by a mere 0.45pp is as effective at raising overall take-up as an attention-boosting intervention that ensures full attention—but only for households in the month after they exit the program. The comparison suggests that under the constraints that policymakers can only intervene current participants, choice nudge is likely a more effective policy.
@@
We test the hypothesis that choice-nudging is more effective than attention-boosting using the Vermont pilot program data. In 2016 and 2017, the Vermont WIC policymakers experimented with two types of text messages in different counties, both of which aimed at increasing the WIC retention rate. The first type of message was attention-boosting. Such messages reminded households who recently dropped out of WIC that they are still likely eligible for the benefits, essentially asking the households to reconsider their choice. The second type of message was choice-nudging. Each month, the WIC policymakers sent out educational messages to inform households of the benefits of continuing WIC participation, improving these households’ perceived utility of the program benefit.
The WIC2Five report concludes that the choice-nudging intervention “did not meet our goal of increasing child retention” by comparing the levels from July 2015 to July 2017. We reanalyze the retention rates data provided by the WIC2Five report by comparing the differences between treated and control sites. Using difference-in-differences (DiD), event study, and permutation test designs, our analysis yields opposing results from the WIC2Five report. Sites that received attention-boosting text messages saw virtually no change in their retention rate; sites that received choice-nudging messages witnessed a significant 8-10pp increase in retention. The results corroborate our counterfactual analysis: a choice-nudging policy is likely to be more effective than an attention-boosting policy when it comes to promoting WIC take-up.
@@
We contribute to three sets of literature. First, the existing works in the welfare take-up literature focus on estimating the causal relationship between observed covariates and the take-up rate (rossin2013wic; chareyron2021does) or the treatment effect of various policy interventions (friedrichsen2018social; shepard2025ordeals). Our work complements this literature by providing a structural framework that disentangles the attention and choice mechanisms behind these causal findings. Second, we propose new identifying conditions that have not been considered by existing inattention model literature (hortaccsu2017power; heiss2021inattention; barseghyan2021discrete; abaluck2021consumers; einav2025selling; agarwal2025demand). Our identifying conditions are tied to the two institutional structures of welfare programs, making our model applicable to many other welfare take-up contexts. We further compare our methodological contribution to the existing inattention model literature in Appendix˜A. Third, we contribute to the long-standing empirical research on WIC (bitler2003wic; rossin2013wic; tiehen2014wic; carlson2021wic; davis2022review). Using different methodologies and datasets, we present two pieces of new evidence that a choice-nudging intervention is likely to be more effective than an attention-boosting intervention in terms of promoting WIC take-up. Further experimental studies to test this hypothesis can be an interesting future research direction.
@@ 2.2 Identification
Following the consideration set literature (abaluck2021consumers; barseghyan2021discrete; agarwal2025demand), we assume that we observe the take-up probability, $P(D_{it}=1\mid D_{it-1},X_{it},Z_{it})$, of a household for any given $\{D_{it-1},X_{it},Z_{it}\}\in\{0,1\}\times\mathcal{X}\times\{0,1\}$. Given the sequential nature of attention and choice, we decompose the observed single-period take-up probability into attention and choice probabilities under Assumption˜2.1 imposed on $\{\mathcal{E}_{it},\Xi_{it}\}$. We will explicitly relax this assumption in the full model.
@@ 3.1 Model setup
Assumption˜2.1 required that the attention utility shock $\mathcal{E}_{it}$ and the choice utility shock $\Xi_{it}$ are i.i.d. over time and mutually independent. The full model relaxes each of these restrictions by allowing for persistent, unobserved household heterogeneity captured by a random effect $Q_{i}$. Intuitively, $Q_{i}$ represents latent time-invariant household traits such as thriftiness (or some part of thriftiness). One component of thriftiness may be explained by observed covariates entering the utility functions, while the orthogonal residual component is unobserved and summarized by $Q_{i}$. In two of the model extensions, we discuss how to relax the orthogonality assumption.
@@ 3.3 Identification
Under Assumption˜3.1, $\beta^{\omega}\in\beta^{\upsilon}\cup\beta^{\chi}$ are constructively identified as
@@
Under Assumption˜3.1, $\beta^{\omega}\in\beta^{\kappa}$ are constructively identified as
@@
Note that all identification results are semiparametric as $\{F,G,H\}$ are not parametrically restricted and are with respect to $X$ variables that have variations both within and across households. For variables that do not vary within or across households, model identifiability relies on the specific parametric assumptions imposed on $\{F,G,H\}$. Nevertheless, the three theorems provide insights into why key structural parameters of the model, such as the influence of household demographics, income, local accessibility on attention and choice, can be learned through a standard panel dataset despite covariates simultaneously entering the attention and choice stages.
Identification of $\{\sigma_{1},\sigma_{2},\sigma_{G},\sigma_{H}\}$
We intuitively explain the identification of $\{\sigma_{1},\sigma_{2},G,H\}$ under parametric assumptions. Assuming that all $\{F,G,H\}$ are normally distributed, we only need to identify the means and variances of the three distributions. Given that we include an intercept term in $X_{it}$, the means of $\{F,G,H\}$ are zero. We only need to identify their variances. Additionally, since $Q_{i}$ is scaled by $\sigma_{1},\sigma_{2}$, we aim to identify $\{\sigma_{1},\sigma_{2},\sigma_{G},\sigma_{H}\}$.
$\sigma_{1}$ and $\sigma_{2}$ are identified from cross-household variation. Assuming that $\sigma_{1},\sigma_{2}>0$. Note that this assumption aligns with reality. A thriftier household is not only more attentive but also has a higher choice tendency. Very heterogeneous and persistent outcomes among households that share similar observed characteristics indicate that household-level heterogeneity has a large variance. If there is a large variance in the waiting time between being eligible and starting participation, then $\sigma_{1}$ is large. If there is a large variance in the participation duration, then $\sigma_{2}$ is large. On the other hand, if there is little difference between take-up behaviors among households with similar observed characteristics, then household-level heterogeneity has a small variance. The identification of $\sigma_{1}$ and $\sigma_{2}$ follows as discussed previously.
$\sigma_{G}$ and $\sigma_{H}$ are identified through within-household comparisons. Our data exhibits high persistence within households along two dimensions: first, households that are already in the program are extremely likely to continue participation; second, households that are not currently enrolled in the program are extremely unlikely to start participation. The first dimension shows that there is little variation in the exogenous part of the choice stage utility shock, $\epsilon_{it}$. In other words, $\sigma_{H}$ is small. The second dimension shows that there is little variation in the exogenous part of the attention stage utility shock, $\xi_{it}$. In other words, $\sigma_{G}$ is small.